Consider a market dominated by a monopolist.

The demand in this market is Q=100-5P. The monopolist faces a constant MC=AC=$4

a. Calculate the monopoly P and Q

b. Calculate the monopoly profit

c. Calculate consumer surplus under monopoly

d. What would P and Q be if this were a perfect competition?

e. What would profit and consumer surplus be if this were a perfect competition?

f. What is the deadweight loss to having the monopoly?

g. If consumers could get together and pay off government officials in order to get rid of the monopoly, what is the max amount would they be willing to pay collectively as a bribe?
Answer #1

Tme=ACESHI 19) Poetstofing Total Revenue = pxQ = (00-04) 98= 2009-es (MR= 20 - 200 Poohit matimization condition MR=MC 20-29-(d) pertect competition pame 20-100 = 4 165 > 19*=807 P= 20 - Do = 20-16